# Bayesian Despeckling of Structured Sources

### v1

@@ 4 Proofs
$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n}\|_{2}^{2}]\geq$ $q_{0}\eta^{2}(2q_{0}(1-{\sqrt{2}\over\Gamma(0.5)})+2q_{0}(1-q_{0})(2-2\Gamma(1.5))+2\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathbbmss{1}_{T\geq 3}])$
$-2M^{2}\Big{(}{(nq_{0})^{-1/4}\over 1-(nq_{0})^{-1/4}}+q_{0}n{\rm e}^{-\lfloor nq_{0}(1+(nq_{0})^{-1\over 4})\rfloor\left(c_{1}^{2}n^{-2\over 3}\over 2n^{-1\over 3}+6\right)}+q_{0}n{\rm e}^{-\lfloor nq_{0}(1-(nq_{0})^{-1\over 4})\rfloor\left(c_{2}^{2}n^{-2\over 3}\over 2n^{-1\over 3}+6\right)}\Big{)},$
where $c_{1},c_{2}\in\mathds{R}^{+}$ are constants. Second line converges to zero as $n\rightarrow\infty$ with $O(n^{-1/4})$. ∎
@@ 5 Numerical experiments
In the following, we first address the practical implementation of the BD-QMAP optimization problem. We then compare the performance of BD-QMAP with the lower bound derived in Theorem 1. Finally, we perform comparative studies on despeckling for piecewise-constant sources and extending the BD-QMAP to be applied on high dimensional images.
5.1 Algorithmic considerations
As noted in Remark 1, addressing the distinct search spaces of the log-likelihood and regularizer in BD-QMAP is essential for practical implementation, and limiting the search to the quantized space ensures feasibility. This restriction enables the discrete suboptimal optimization to be efficiently performed using the Viterbi algorithm [32]. Therefore, to tackle the optimization problem of BD-QMAP, we begin by limiting the search space in (3) to $b$-bit quantized realizations, denoted as $\mathcal{X}^{n}_{b}$. We call this suboptimal despeckler BD-QMAPb, as it is optimizing the both terms of (3) in the $b$-bit quantized space. In other words, BD-QMAPb solves the following optimization
$\operatorname*{arg\,min}_{u^{n}\in\mathcal{X}^{n}_{b}}{1\over n}\sum_{i=k}^{n}(\log u_{i}^{2}+\frac{y_{i}^{2}}{u_{i}^{2}}+\frac{\lambda}{b}w_{u_{i-k+1}^{i}}),$ (33)
where $c_{\mathbf{w}}(u^{n})$ is scaled by $(n-k+1)/n$ without an impact on the optimal solution as it can be absorbed in the hyper-parameter $\lambda$. Since similar to the the additive case, the negative log likelihood function of $u^{n}$ is the sum of $n$ element-wise distinct functions, with the $i$-th function only depending on $u_{i}$, the Viterbi algorithm [32] can still be employed to solve (33).
The Viterbi algorithm operates on a trellis diagram with $n$ stages, each having $|\mathcal{X}_{b}|^{k}$ states. State $s_{i}$ in the diagram corresponds to the subsequence $u_{i-k+1}^{i}$ of length $k$ in the candidate solution. Let the cost associated to the state $s_{i}=u^{i}_{i-k+1}$ in the diagram follows the recursive rule as,
$C_{i}(s_{i})=\min_{s^{\prime}\in\mathcal{S}(s_{i})}(\log u_{i}^{2}+\frac{y_{i}^{2}}{u_{i}^{2}}+\frac{\lambda}{b}w_{u^{i}_{i-k+1}})+C_{i-1}(s^{\prime}),$
for any $i\in\{k,\dots,n\}$, where $\mathcal{S}(s_{i})$ denotes the set of all closest states that transition into the state $s_{i}$ and $|\mathcal{S}(s_{i})|=|\mathcal{X}_{b}|$. The state $s^{\prime}=u_{i-k}^{i-1}$ represents the previous state leading to $s_{i}$. Thus BD-QMAPb solution can be found recursively starting minimizing $C_{n}(s_{n})$.
@@ 5.2 Theoretical lower bound
In Theorem 1, a lower bound on the best achievable reconstruction error was derived for the piecewise constant source. To explore this bound in practice, we compare it to an ideal despeckler for piecewise constant sources, referred to as the genie-aided maximum likelihood (ML) despeckler. This estimator has access to the exact jump locations in the ground truth data in addition to the noisy observations. With such side information, the reconstruction error is minimized using Equation 8 for each constant segment between consecutive jump locations. For comparison, we evaluated the performance of this genie-aided despeckler against BD-QMAP. The BD-QMAP implementation, as detailed in Section 5.3.1, was applied to piecewise constant sources with $q_{0}\in\{0.001,0.01,0.05,0.1\}$. Each data point in the figure represents the average error over a test set of 100 realizations with fixed $q_{0}$, each of length $n=10^{6}$.
@@ 5.3 Comparative analysis of despeckling for piecewise constant sources
For numerical experiments with the piecewise constant source introduced in Section 3.2 we assume the absolute continous distribution to be uniform on the interval $[0,1]$. A sample realization of this source under speckle noise is shown in Figure 2. Considering this source, the pivotal parameter which directly affects the performance of the despeckler is $q_{0}$, i.e., the probability of signal level-change which also determines the level of structure for the analog source [29]. To learn the weights of the Equation 2 to be used in the regularizer term of the BD-QMAP loss function, for a fixed $q_{0}$, we sample the random process for $10^{7}$ points and find the empirical second order distribution by counting the occurances of the $b$ bit quantized samples of length two. Key advantages of BD-QMAP over other despecklers in the literature in practice are two fold. First, the BD-QMAP assumes no upperbound on the speckle power level on the observation as opposed to [13, 14], and also [8] in which the filter is bypassed completely in high noise level regimes. Second, the BD-QMAP method is applicable to any given source without requiring restrictive assumptions on the signal’s prior distribution. This flexibility contrasts with the strong assumptions imposed by other Bayesian filters, such as those in [7, 10, 6]. Each entry in the Table 1 is the despeckling performance for a test dataset of size $100$ of piecewise constant signals of length $n=10^{6}$.
@@ 5.3.1 BD-QMAP implementation
Using the Viterbi implementation described in Section 5.1 for $k=2$, we can efficiently implement BD-QMAPb. As noted in Remark 1, the fidelity term in BD-QMAP optimization is not restricted to a quantized space. Therefore, to enhance the BD-QMAPb solution, we refine it by identifying the detected jump locations and computing the ML estimates on the intervals between them using Equation 8. This refinement significantly improves performance, especially for highly structured sources (i.e., when $q_{0}$ is small). As shown in Table 1, the despeckling performance of BD-QMAPb for highly structured sources improves substantially when the initial estimate is refined for the final BD-QMAP solution. Notably, even a low-complexity implementation with $b=2$ is effective in detecting jump locations, with the refinement step yielding significant gains.
@@ 5.3.2 Regularization and quantization in BD-QMAP
To explore the effectiveness of the regularizer of the BD-QMAP in finding the structure of the source, and yeilding lower errors, we conducted experiments for a range of values for $\lambda$ in Equation 3, as shown in Figure 3. It can be seen that when $\lambda=0$ the despeckler is not able to distinguish between different levels of structuredness (as measured by $q_{0}$), however a small regularization would let the denoiser to recover the empirical distribution of the ground-truth distribution in the reconstruction. Finally, as expected, a large value of $\lambda$ would decrease the fidelity of the reconstruction to the observed noisy signal. The number of bits used to quantize the source for the learned weights significantly impacts both the final performance and the time complexity of the optimization problem in Equation 3, as the size of the search space increases exponentially with the number of bits.
In general, achieving better reconstruction results necessitates selecting the number of bits $b$ in BD-QMAP as high as computationally feasible. However, for highly structured data (e.g., high-resolution images), a relatively small number of bits can suffice to uncover the structure of the source from noisy observations. As demonstrated in Figure 3, when $q_{0}=0.001$ (indicating a highly structured source), the suboptimal solution of BD-QMAPb with just 2 bits effectively captures the source structure. Upon refining the solution to account for quantization error, it achieves a higher PSNR than BD-QMAPb with 3 bits. This experiment suggests that for highly structured sources, selecting fewer bits is often sufficient to match the distribution, avoiding the computational challenges associated with higher bit counts.
@@ 5.3.3 Linear adaptive filters
All the linear adaptive filters in Table 1 require a window size hyper-parameter that defines the area over which the local statistics will be calculated. In these experiments this value is set to $1/2q_{0}$, half of the expected interval lengths of constant pieces in the source.
We also conduct experiments with an Enhanced version of the filters which is based on the idea of thresholding local variance, as suggested in [8]. To improve the final performance of the filters, three statiscical classes are considered which are known as heterogeneity adjustments in the literature. The simple modification is to distinguish between three region of operation: homogeneous, heterogenous, and strongly speckled. To decide in which region the filter operates, if the observation variation coefficient defined as $C_{Y}\triangleq\frac{\operatorname{Var}(Y)}{\operatorname{E}[Y]}$ of the local window falls below a threshold of noise variation coefficient, similarly defined to be $C_{W}\triangleq\frac{\operatorname{Var}(W)}{\operatorname{E}[W]}$, we only do keep the mean window mean value as the despeckled pixel. On the other hand, if $C_{Y}\geq C_{\max}$, the filter output is bypassed and the despeckled pixel is equal to the speckled observed pixel, where empirically the optimal value of $C_{\max}$ is set to $\sqrt{3}C_{W}$ [8]. The Enhanced version of the filters mentioned in Table 1 refer to applying the filter only when $C_{W}\leq C_{Y}\leq C_{\max}$.
@@ 6 Conclusion
A novel Bayesian despeckling method is proposed for structured sources. Its performance is analyzed for piecewise constant structured sources, and a lower bound on the minimum achievable MSE is derived. The results demonstrate that the despeckler effectively detects edges under speckle noise, achieving superior performance compared to other theoretically grounded approaches.

### v2

@@ 4.3 Proof of Corollary 1
$\operatorname{E}[{1\over n}\|X^{n}-{\hat{X}}^{n}\|_{2}^{2}]\geq$ $2q_{0}\eta^{2}\Big{(}q_{0}c_{1}+q_{0}^{2}c_{2}+\operatorname{E}[T(1-{1\over\sqrt{{\rm e}}}{(1-{1\over T})^{T\over 2}\over(1-{2\over T})^{T-1\over 2}})\mathds{1}_{T\geq 3}]\Big{)}$
$-2x_{M}^{2}\Big{(}{(nq_{0})^{-1/4}\over 1-(nq_{0})^{-1/4}}+q_{0}n{\rm e}^{-\lfloor nq_{0}(1+(nq_{0})^{-1/4})\rfloor\left(d_{1}^{2}n^{-2/3}\over 2n^{-1/3}+6\right)}+q_{0}n{\rm e}^{-\lfloor nq_{0}(1-(nq_{0})^{-1/4})\rfloor\left(d_{2}^{2}n^{-2/3}\over 2n^{-1/3}+6\right)}\Big{)},$
where $d_{1},d_{2}\in\mathds{R}^{+}$ are constants. Second line converges to zero as $n\rightarrow\infty$ with $O(n^{-1/4})$.
@@ 5 Numerical Experiments
This section evaluates the performance of the proposed BD-QMAP method. We begin by addressing the practical implementation challenges of the BD-QMAP optimization problem. Next, we compare its performance against the theoretical lower bound derived in Theorem 1. Finally, we investigate BD-QMAP’s application to piecewise-constant sources.
5.1 Implementation of BD-QMAP
[…]
$\operatorname*{arg\,min}_{u^{n}\in\mathcal{X}^{n}_{b}}\frac{1}{n}\sum_{i=k}^{n}\left(\log u_{i}^{2}+\frac{y_{i}^{2}}{u_{i}^{2}}+\frac{\lambda}{b}w_{u_{i-k+1}^{i}}\right),$
where the term $c_{\mathbf{w}}(u^{n})$ is scaled by $(n-k+1)/n$. This scaling has no impact on the optimal solution, as it can be absorbed into the hyperparameter $\lambda$.
[…]
The Viterbi algorithm operates on a trellis diagram with $n$ stages, each containing $|\mathcal{X}_{b}|^{k}$ states. A state $s_{i}$ in the diagram corresponds to the subsequence $u_{i-k+1}^{i}$ of length $k$ in the candidate solution. The cost associated with state $s_{i}=u^{i}_{i-k+1}$ is updated recursively as:
$C_{i}(s_{i})=\min_{s^{\prime}\in\mathcal{S}(s_{i})}\left(\log u_{i}^{2}+\frac{y_{i}^{2}}{u_{i}^{2}}+\frac{\lambda}{b}w_{u^{i}_{i-k+1}}+C_{i-1}(s^{\prime})\right),$
for $i\in\{k,\dots,n\}$, where $\mathcal{S}(s_{i})$ denotes the set of all states that transition into $s_{i}$, and $|\mathcal{S}(s_{i})|=|\mathcal{X}_{b}|$. The state $s^{\prime}=u_{i-k}^{i-1}$ represents the preceding state leading to $s_{i}$. The BD-QMAPb solution is obtained by minimizing $C_{n}(s_{n})$ recursively.
@@ 5.2 Theoretical Lower Bound
In Theorem 1, we derived a theoretical lower bound on the minimum achievable reconstruction error for piecewise-constant sources. To explore this bound empirically, we compare it with the performance of an ideal despeckler referred to as the genie-aided maximum likelihood (ML) despeckler. This ideal estimator assumes access to the exact jump locations in the ground truth signal in addition to the noisy observations. With this side information, the reconstruction error is minimized by applying Equation 8 to each constant segment between consecutive jump locations. We evaluate this estimator and compare it to the lower bound, along with the experiments on piecewise constant source despeckling using BD-QMAP, as explained in the next section.
@@ 5.3 Comparative analysis of despeckling for piecewise constant sources
This subsection evaluates the effectiveness of BD-QMAP in despeckling piecewise-constant sources. The results are compared against several well-known despeckling methods, including linear adaptive filters and enhanced versions thereof. Additionally, the role of regularization and quantization in BD-QMAP is explored to understand its impact on reconstruction quality and computational efficiency.
For numerical experiments with the piecewise-constant source introduced in Section 3.2, we assume the absolutely continuous distribution is uniform over the interval $[0,1]$. A sample realization of this source under speckle noise is shown in Figure 1. For this source, the pivotal parameter affecting despeckler’s performance is $q_{0}$, the probability of a signal level change, which determines the degree of structure in the source [29]. To train the weights of Equation 2 used in the BD-QMAP regularizer, we sample the random process for $10^{7}$ points with fixed $q_{0}$ and compute the empirical second-order distribution by counting occurrences of $b$-bit quantized sample pairs.
The BD-QMAP method offers two key practical advantages over existing despecklers. First, unlike methods such as [13, 14] and [8], BD-QMAP does not assume an upper bound on speckle power levels in the observations, making it robust to high-noise regimes where other filters fail or bypass the input. Second, BD-QMAP is flexible and does not require restrictive assumptions on the signal’s prior distribution, in contrast to Bayesian filters like those in [7, 10, 6]. These features make BD-QMAP well-suited to a wide range of structured sources. Table 1 summarizes the despeckling performance for test datasets of size 100, each containing piecewise-constant signals of length $n=10^{5}$.
@@ 5.3.1 BD-QMAP implementation
Using the Viterbi implementation described in Section 5.1 for $k=2$, we efficiently implement BD-QMAPb. As noted in Remark 1, the fidelity term in BD-QMAP is not restricted to the quantized space. Therefore, we refine the BD-QMAPb solution by identifying detected jump locations and applying ML estimation on the intervals between jumps using Equation 8. This updated estimation is called BD-QMAP in our experiments. This refinement significantly improves performance, particularly for highly structured sources (i.e., small $q_{0}$). As shown in Table 1, even a low-complexity implementation with $b=2$ effectively detects jump locations, and the refinement step yields substantial performance gains. In Figure 2 the performance of BD-QMAPb and its refined version, BD-QMAP, is compared to the theoretical lower bound from Theorem 1 and the genie-aided ML estimator from Section 5.2.
@@ 5.3.2 Regularization and quantization in BD-QMAP
To evaluate the impact of the BD-QMAP regularizer, we conduct experiments across a range of values for the hyperparameter $\lambda$ in Equation 3, as shown in Figure 3. When $\lambda=0$, the despeckler fails to distinguish between sources with different levels of structure (as measured by $q_{0}$). Introducing a small regularization allows the denoiser to recover the empirical distribution of the ground truth in the reconstruction. However, excessively large values of $\lambda$ reduce fidelity to the noisy signal.
The number of bits used for quantization ($b$) also significantly impacts performance and computational complexity. While larger $b$ generally yields better reconstructions, for highly structured sources (e.g., high-resolution images), even small $b$ can effectively capture source structure. As demonstrated in Figure 3, for $q_{0}=0.001$, the suboptimal BD-QMAPb solution with just $b=2$ bits captures the source structure well. After refinement, it achieves higher PSNR than BD-QMAPb with $b=3$, suggesting that smaller $b$ can be sufficient for highly structured data while reducing computational costs.
@@ 5.3.3 Linear adaptive filters
All linear adaptive filters in Table 1 require a window-size hyperparameter to define the region over which local statistics are computed. In our experiments, this value is set to $1/2q_{0}$, corresponding to half the expected interval length of constant signal pieces. This choice ensures the filters adapt to the level of structure in the source.
We also experiment with enhanced versions of these filters, incorporating heterogeneity adjustments as proposed in [8]. This approach classifies regions as homogeneous, heterogeneous, or strongly speckled based on the local coefficient of variation of the observations, defined as $C_{Y}=\operatorname{Var}(Y)/\operatorname{E}[Y]$.
The enhanced filter operates as follows: If $C_{Y}$ falls below the noise coefficient of variation $C_{W}=\operatorname{Var}(W)/\operatorname{E}[W]$, the filter outputs the mean value of the local window. If $C_{Y}\geq C_{\max}$, where $C_{\max}=\sqrt{3}C_{W}$, the filter is bypassed, and the observed pixel is retained. Otherwise, the filter is applied normally. The enhanced versions of the filters in Table 1 incorporate these modifications to improve despeckling performance.
@@ 6 Conclusion
A novel Bayesian despeckling method, BD-QMAP, is proposed for structured sources, providing a theoretically grounded approach to addressing speckle noise. Its performance is analyzed for piecewise-constant structured sources, with a derived lower bound on the minimum achievable MSE. Experimental results highlight the effectiveness of BD-QMAP in despeckling piecewise-constant sources, achieving state-of-the-art performance.
However, the theoretical understanding of despeckling and its fundamental limits remains significantly less developed compared to denoising in additive noise models, presenting substantial opportunities for further research. Additionally, extending the application of QMAP-based methods to image despeckling is a promising direction, which we plan to explore in future work.